Water heater is rated 1200w, 220v. Calculate

(i) it's resistance

(ii) the cost of operating 5 water heaters for 5 hours.(1kwh)=$34.00


1
Expert's answer
2021-06-11T11:22:25-0400

(i) The resistance is


"R=\frac{V^2}{P}=40.3\Omega."

(ii) 5 heaters consume 5·1200=6000 W, or 6 kW. 1 kW costs $34, 6 kW cost 6·$34=$204.